OpenStudy (anonymous):

WILL GIVE MEDAL The distance traveled by an object can be modeled by the equation d = ut + 0.5at2 where d = distance, u = initial velocity, t = time, and a = acceleration. Solve this formula for a. Show all steps in your work.

OpenStudy (anonymous):

d=ut+0.5at2 subtract ut from both sides d−ut=0.5at2 divide both sides by 0.5 (this is the same as multiplying by 2) 2(d−ut)=at2 divide both sides by t^2 2d−2utt2=a

OpenStudy (anonymous):

can you show the work? like 2x+3y=1,200 -2x -2x

OpenStudy (anonymous):

im having trouble understanding. @ramrodz31